The concentration of a 100ml solution containing X grams of Na2Co3 is Y M.the values of X and Y are..?

Grade:11

1 Answers

Gaurav
8 years ago
Molar mass of Na2CO3 = 106 g/mol
Number of Moles of x gram Na2CO3 = X/106
Volume of Solution = 0.1 L
Molarity = (x/106)/0.1

or
y =(x/106)/0.1
or
10.6y = x